OpenStudy (anonymous):

Find f(6) if f(x) = x^2 ÷ 3 + x. 4 10 18

OpenStudy (anonymous):

f(6) if f(x) = x^2 ÷ 3 + x. = (6)^2 / 3+6=

OpenStudy (anonymous):

=36/3+6 =12+6=18

OpenStudy (anonymous):

thank you:)